OpenStudy (anonymous):

Can some one show me how to get the derivative of sin2x using the limit definition?

OpenStudy (ash2326):

Derivative is defined as \[f'(x)=lim_{h\to 0} \frac{f(x+h)-f(x)}{x+h-h}\] here f(x) = sin 2x \[f'(x)=lim_{h\to 0} \frac{ \sin 2(x+h) - sin 2x}{h}\]Let's expand sin (2x+2h)= sin 2x cos 2h+ sin 2h cos 2x \[f'(x) =lim_{h\to 0} \frac{ \sin 2x \cos 2h+ \sin 2h cos 2x- \sin 2x}{h}\] Now cos 2h -->1 as h--->0 so \[f'(x) =lim_{h\to 0} \frac{ \sin 2x+ \sin 2h \times cos 2x- \sin 2x}{h}\] \[f'(x) =lim_{h\to 0} \frac{ \sin 2h \times cos 2x}{h}\] We know \[\lim_{h \to 0} \frac{sin h}{h}=1\] so \[f'(x) =lim_{h\to 0} \frac{2 \times \sin 2h \times cos 2x}{2h}\] we get \[f'(x)= 2cos 2x \lim_{h \to 0} \frac{sin 2h}{2h}\] Finally \[f'(x)= 2 \times \cos 2x \times 1\]

OpenStudy (anonymous):

Thanks : )